What is the difference between target profit analysis and break-even analysis? Give one example of each
Answer #1

Target profite analysis methos is a kind of profit volume analysis. The main objective of target profit analysis is to achieve a sales volume that can cover the fixed and the variable costs and can achieve a targeted profit for a time period. The target profit indicates the necessary sales revenue to be earned to achieve the targeted profit.

For example, X company aspire to earn a profit of 250,000 USD in the coming quarter. Fixed cost expected to stand 700,000 USD and the variable cost per unit is 500 USD , while sales price per unit is 900 USD.

Target sales will be= (Fixed cost+targeted profit)/(sales price-variable cost)

=(700,000-250,000)/(900-500)

=450,000/400=

2375 units

Break even analysis refers to a point where the total costs and total revenue gets equal. The break even analysis is generally conputed to determine the number of units or revenue which is required to cover the total costs.

For example, fixed costs of Y is 100,000 USD and variable cost is 2 USD per unit. Selling price of the product is 12 USD.

Break even sales will be= fixed cost/(sales price-variable cost)

=100,000/12-2

=10,000 units